### v1

@@ Abstract
Introduction Augmented Reality (AR) has the potential to change the way we engage in knowledge work. Commercial off-the-shelf devices, such as the Apple Vision Pro [3] and Lenovo ThinkReality Glasses [55], are increasingly promoted as productivity tools. Previous research has also highlighted several benefits that AR offers in this context, such as reducing visual distractions and allowing the personalization of shared workspaces [54]. Furthermore, AR has been shown to support the in situ display of digital information [18] and facilitate collaborative embodied brainstorming and sensemaking [62].
@@
A particularly promising affordance of AR for knowledge work is its ability to present information and applications without being limited to physical monitors [70]. Using a head-mounted display (HMD), virtual windows, icons, and menus that were previously confined to computer screens can be seamlessly integrated into the user’s physical environment, allowing them to be displayed anytime and anywhere, with variable size and appearance. This support for an effectively infinite display space provides significant productivity benefits over conventional devices by minimizing context-switching costs and improving user awareness of peripheral information [63]. Additionally, these properties may be particularly beneficial in mobile scenarios, facilitating work in diverse contexts, such as public or confined spaces where the size of physical hardware is limited [11].
@@
Many companies and researchers suggest AR as a medium will eventually replace computers and phones in the “office of the future” [40], enabling all-day knowledge work on the go. However, prior studies evaluating AR’s support of knowledge work have predominantly been conducted in laboratory settings and over relatively short timeframes (e.g., [11, 70]). Although these earlier studies are valuable for validating the feasibility of performing knowledge work with current AR devices and informing future developments, they may not fully reflect user experiences during everyday computing activities [1]. Furthermore, the limited duration of these studies may obscure our understanding of how this technology shapes people’s behaviors and attitudes once they have had time to adjust to its novelty [7]. Therefore, we see it as beneficial to explore how people use AR for work in their daily environments over time.
[…]
@@
Our analysis of the collected survey data, workspace photographs, and interviews provides insights into how considerations around task, environment, social, and physical comfort influenced participants’ usage behaviors, including the number, size, and arrangement of windows within their virtual workspaces. We also report on instances where participants naturally used the device to complement physical displays and tasks. Finally, our analysis highlights the current values and challenges of using an AR laptop for tasks in everyday work settings. In our discussion, we reflect on how our findings, derived from a more ecologically-valid and longitudinal approach to investigating AR usage, relate to prior empirical studies. We also highlight implications for future design and research, including directions for developing interaction techniques, onboarding approaches, and adaptive systems.
@@
  • A detailed analysis of AR usage patterns from a longitudinal diary study conducted in situ within people’s daily work environments. We report on usage considerations, virtual display configuration patterns, and hybrid usage involving both physical displays and tasks
@@ 1 Related Work
Our work builds upon prior Mixed Reality (MR) research, including productivity-oriented applications, window-based interfaces, and in-the-wild as well as longitudinal studies.
@@ 1.1 Productivity in Mixed Reality
In prior research, there is a significant body of work leveraging devices across the MR continuum, including Augmented and Virtual Reality (AR/VR), for productivity-related applications. Many systems and prototypes have been developed to support tasks such as sensemaking [59], working with spreadsheets [37], data visualization [16], and collaboration [74]. These systems use a variety of novel interaction techniques, ranging from multi-device input for navigating information [53, 13] to gaze-based approaches for accessing applications [60].
Beyond promising benefits for application-specific tasks, for many commercial stakeholders and researchers, MR represents a future pervasive medium for work, much like how computers and mobile phones are today [40, 39, 35]. Numerous studies have highlighted the advantages that MR offers as a ubiquitous tool for work. For instance, MR’s ability to personalize a user’s visual work environment, such as replacing an open office background with a calming natural scene, has been shown to reduce distractions and stress while inducing a state of flow [79, 54, 92]. Schneider et al. [81] and Grubert et al. [40] examined the potential privacy benefits of the HMD form factor, which restricts the visibility of content to the wearer. Last but not least, MR relieves information displays of physical limitations [12, 40]. Using an HMD to present virtual content, including conventional windows, icons, and menus, MR provides more flexibility in where and how information is displayed [70]. As promoted by providers like Apple, Meta, and Sightful, it effectively provides “boundless” workspace [86]. This expanded screen real-estate may facilitate complex multi-window operations by serving as a type of external memory [72, 2, 9], enable more efficient multitasking [28], provide quicker peripheral access to information [60, 29, 41], and enhance immersion [31].
@@ 1.2 Window-based Mixed Reality
MR systems are often linked to novel interaction techniques, but there has also been a persistent interest in using these devices to enhance knowledge work by expanding traditional 2D displays and windows beyond the limitations of physical screens and into the third dimension [72]. Early work, such as Feiner et al. [34], explored various methods to register windows, including head-fixed and world-fixed approaches, as well as how to create connections between windows and physical objects. Raskar et al. [77] investigated the possibilities of expanding an office space with AR using projectors and cameras. Many recent commercial MR workspaces, such as the Varjo Workspace [93], Immersed [46], and vSpatial [94], now enable streaming of desktop elements into MR as virtual windows and support conventional input devices (i.e., mouse and keyboard) for interaction. Using conventional 2D displays and desktop input devices for MR can ease the transition to virtual displays for users coming from personal computers, while also providing access to familiar and highly refined interfaces and applications [72]. Therefore, in our study, we focus on emergent user behaviors when using an AR laptop that features a window-based immersive display along with keyboard and touchpad input.
Several prior studies have investigated the usage of window-based MR. For example, Ens et al. [32] devised a design framework to classify window-based mixed reality workspaces based on windows in seven dimensions, such as whether the windows used an ego- or exocentric reference frame or were movable. Several prior studies [19, 70, 58] observed that users preferred to center main task content while using peripheral areas for supporting information. Su and Bailey [90] recommended against placing virtual contents behind the user. McGill et al. [63] reported on several common window configurations for general productivity tasks, including vertical, horizontal, two-plus-two, and following a “personal cockpit” [33]. As immersive technologies become more pervasively used, the context of use also becomes a critical consideration [39]. For example, the physical environment can hinder or enhance interactions [20]. Similarly, social acceptability has been shown to play an important role in the use of virtual monitors in airplanes [67] and other shared transit modalities [64]. While many usage patterns have been observed in shorter laboratory studies, our research investigates how these patterns may manifest in more realistic environments and over extended periods.
@@ 1.3 In-the-Wild and Longitudinal Studies of MR
In the current literature, there is a growing number of in-the-wild and extended-use studies of MR systems. In-the-wild studies of MR have demonstrated effectiveness and data quality across various application areas, including education [75], accessibility [80], recreational sports [24, 23], and military training [52]. Researchers have also investigated the long-term effects of MR usage on depth perception [51], social interactions [44, 8], and within industrial scenarios [38]. The most relevant to our work are previous studies on the use of MR for everyday productivity tasks.
In early investigations, many researchers engaged in informal self-experimentation. Starner [89] and Mann [15] independently designed and wore AR headsets almost daily for over a decade. During this time, they reported using the device for tasks such as checking the Internet and taking notes during face-to-face meetings [89], along with experiences of backlash and more negative physiological aftereffects [15]. Recent studies have continued this line of work with more participants, often with the aim of quantifying the effects. For example, Steinicke and Bruder [88] and Nordahl et al. [68] examined the impact of prolonged VR usage on key usability factors, including simulator sickness, over periods of 24 hours and 12 hours, respectively. Lu et al. [61] evaluated a glanceable AR prototype with three participants over three days. Guo et al. [42] and Shen et al. [82] report on a study in which 27 participants worked eight hours each in virtual and physical office environments, focusing on emotional and physiological needs, as well as visual and physical discomfort. Biener et al. [12, 10] compared VR work with a regular physical environment over five days with 16 participants. Pavanatto et al. [70] recently examined user and bystander experiences of extended reality (XR) usage in public, studying how users perform their own tasks in XR at various campus locations (e.g., library, dining).
@@ 2 Study
In the following, we first provide a detailed description of our apparatus and the rationale behind its use. We then report on our study and analysis methodology.
@@ 2.1 Apparatus
After pilot testing several devices and approaches, we chose to use the Sightful Spacetop EA (Early Access) [85], shown in Figure 1. The device consists of a pair of optical see-through display glasses (Xreal Light, $85$ grams, OLED display panels, $1920\times 1080$ pixels per eye, $72$ Hz refresh rate, ${}52^{\circ}$ diagonal field of view) for output attached to a standard keyboard and trackpad for input. It has a Qualcomm Snapdragon XR2 processor, a Kryo 585TM 8-core 64-bit CPU, an AdrenoTM 650 GPU, and 8GB of LPDDR5 RAM. The device runs the Spacetop OS [84], which enables users to flexibly open, close, resize, and move one or more window-based web applications (e.g., Google workspace application, WhatsApp, ChatGPT) on top of a virtual “spatial canvas” [87] — a unified display space that is curved over a cylinder surrounding the user, similar to the concept described by Pavanatto et al. [72]. The spatial canvas is notably world-anchored [34] and can be zoomed in and out of, panned from side to side, height adjusted, tilted, and re-centered over the device keyboard [83]. We chose the device over video passthrough options like the Meta Quest 3 [65] or Apple Vision Pro [3] mainly because of its lighter and more ergonomic glasses form factor, which is essential for prolonged productive tasks [12].
@@ 2.2 Experimental Design and Procedure
Our particular approach was guided by several studies of methodological similarity (e.g., [26, 14, 47]). Diary studies are a common qualitative research method in HCI [25], widely used to understand behavior over longer periods, such as working within multi-device ecologies [47] or work-from-home conditions during the COVID-19 pandemic [14]. In the following, we describe how we adapted these methodologies to capture in-depth data on knowledge workers’ use and perceptions of AR in their daily activities. The study was approved by the IRB of the participating research institutions.
@@
We deployed the Sightful EA device to each participant for two weeks (14 days). During this usage period, participants were instructed to complete a minimum of 10 sessions, with each session involving at least 30 minutes of continuous use. Participants had the option to complete additional sessions for extra compensation.
Participants were encouraged to use the device during their daily activities and tasks, which they would typically complete regardless of their participation in the study, in the familiar locations they frequented as part of their routines. We deliberately chose not to set specific constraints on the task and environment, in order to gather more ecologically valid data on participants’ usage behaviors. We set individual sessions at a minimum of 30 minutes to ensure that participants had enough time to engage in meaningful tasks. We evaluated usage across ten daily sessions to allow participants to acclimate to the device and understand how their usage changes with different tasks and extended exposure, without overburdening them.
@@
We conducted pilot tests of the survey with four knowledge workers, who were recruited through snowball sampling. These pilot sessions helped us identify core questions, clarify wording, and ensure data quality. We revised the survey after each round of feedback. The results from these pilot sessions were not included in our analysis.
@@
Participants received \$$\$150$ for completing ten usage sessions and the exit interview. They received \$$\$15$ for each additional session they completed beyond the minimum requirements. Usage sessions took place from March 19, 2024, to August 29, 2024. We deployed between two and three devices simultaneously to speed up data collection.
@@ 2.3 Participants
We recruited 14 people (aged 22 to 56 years, $M$ = 30, $SD$ = 11; 9 male, 5 female) with normal or corrected vision to participate in the study. We recruited university students, researchers, and staff for our study due to their diverse tasks, extensive computer use in their work, and availability [6]. Participants self-reported low to moderate experience with AR ($M$ = 2.6, $SD$ = 1.8) and VR ($M$ = 3.1, $SD$ = 1.7) on a scale from 1 (low) to 7 (high). Participants indicated that they regularly use 2 to 7 ($M$ = 4, $SD$ = 1) digital devices (i.e., mobile phones, laptops, tablets, and smartwatches). They reported spending approximately 5 to 14 hours each day ($M$ = 7, $SD$ = 2) on their primary device, which was either a laptop or desktop computer. All participants reported working remotely to some extent, typically at least once a week. They also reported having normal or corrected-to-normal vision.
@@ 2.4 Data Collection and Analysis
In addition, participants collectively submitted 140 photographs of their virtual workspaces and 109 photographs of their physical environments. All interviews with participants were transcribed.
Figure 2: Themes — We report on how considerations, such as the participants’ tasks, influenced their workspace arrangement behaviors, such as the number of windows they opened. In addition, we highlight alternative workflows that involve the use of AR in tandem with physical displays and tasks. Finally, we summarize participants’ perceptions of the value and challenges of using AR.
@@
The final codebook comprised 171 individual codes that detailed the tasks participants performed using the AR laptop, the physical environments in which they worked, how they organized their virtual workspace for their tasks, and their perceptions regarding the value and limitations of the device. Using this shared codebook, two authors collaboratively conducted a final round of coding on the original data. They applied the codes from the codebook to each instance, labeling every survey entry and interview with zero or more codes, based on a unanimous decision. We do not report inter-rater reliability because following best practices from Hammer and Berland [43], coding for each instance was unanimously agreed upon by annotators. The same two authors also led the clustering process, organizing the codes into themes. These themes were refined iteratively with feedback from the remaining members of the research team, resulting in the findings presented below.

### v2

@@ 1 Introduction
Augmented Reality (AR) has the potential to change the way we engage in knowledge work. Commercial off-the-shelf devices, such as the Apple Vision Pro [3] and Lenovo ThinkReality Glasses [52], are increasingly promoted as productivity tools. Previous research has also highlighted several benefits that AR offers in this context, such as reducing visual distractions and allowing the personalization of shared workspaces [51]. Furthermore, AR has been shown to support the in situ display of digital information [17] and facilitate collaborative embodied brainstorming and sensemaking [61].
@@
A particularly promising affordance of AR for knowledge work is its ability to present information without being limited to physical monitors [70]. Using a head-mounted display (HMD), application contents that were previously confined to computer screens can be seamlessly integrated into the user’s physical environment, allowing them to be displayed anytime and anywhere, with variable size and appearance. This support for an effectively infinite display space provides significant productivity benefits over conventional devices by minimizing context-switching costs and improving user awareness of peripheral information [62]. These properties are especially useful in mobile scenarios, enabling work in public or confined spaces [10].
@@
Many companies and researchers suggest that AR will eventually replace computers and phones in the “office of the future” [37], enabling all-day knowledge work on the go. However, prior studies evaluating AR’s support of knowledge work have predominantly been conducted in laboratory settings and over relatively short timeframes (e.g., [10, 70]). Although these earlier studies are valuable for validating the feasibility of performing knowledge work with current AR devices and informing future developments, they may not fully reflect user experiences during everyday computing activities [1]. Furthermore, the limited duration of these studies obscures our understanding of how this technology shapes people’s behaviors once they have had time to adjust to its novelty [7]. Therefore, we see it as beneficial to explore how people use AR for work in their daily environments over time.
In this paper, we present the results of a diary study investigating participants’ usage of an AR laptop — Sightful’s Spacetop EA (Early Access) — for daily work tasks and activities. The Spacetop EA is an optical see-through AR HMD connected to a laptop base with standard keyboard and trackpad for input (Section 3.1). The device enables people to work on real productive tasks in their own physical environments. Notably, it supports flexibly opening and arranging virtual windows for common applications used in knowledge work, similar to the interfaces explored by Pavanatto et al. [72]. In our study, we deployed this AR laptop for 14 participants to use during their workday over the course of two weeks in 40-minute daily sessions. During the study, we collected photographs of their virtual workspace and physical environment, along with post-session survey responses about their usage and perception of the device. Participants collectively completed a total of 103 hours of work using the AR laptop ($\scriptstyle\sim$7 hours per participant) across 143 sessions ($\scriptstyle\sim$10 sessions per participant). The study concluded with $\scriptstyle\sim$30-minute interviews, encouraging participants to reflect more deeply on their overall experience.
@@
Our analysis of the survey data, workspace photographs, and interviews provides insights into how considerations around task, environment, social, and physical comfort influenced participants’ usage behaviors, including the number, size, and arrangement of windows within their virtual workspaces. We also report on instances where participants naturally used the device to complement physical displays and tasks. Finally, our analysis highlights the current values and challenges of using an AR laptop for tasks in everyday work settings. In our discussion, we reflect on how our findings, derived from a more ecologically-valid and longitudinal approach to investigating AR usage, relate to prior empirical studies. We also highlight implications for future design and research, including directions for developing interaction techniques, onboarding approaches, and adaptive systems.
@@
  • An analysis of AR usage patterns from a longitudinal diary study conducted in situ within people’s daily work environments. We report on usage considerations, virtual display configuration patterns, and hybrid usage involving both physical displays and tasks
@@ 2.1 Productivity in Mixed Reality
In prior research, there is a significant body of work leveraging devices across the Mixed Reality (MR) continuum, including Augmented and Virtual Reality (AR/VR), for productivity-related applications. Many systems and prototypes have been developed to support tasks such as sensemaking [56], working with spreadsheets [34], and collaboration [74]. These systems use a variety of interaction techniques, ranging from multi-device input for navigating information [50, 12] to gaze-based approaches for accessing applications [58].
For many commercial stakeholders and researchers, MR represents a future pervasive medium for work, much like how computers and mobile phones are today [37, 36, 31]. Numerous studies have highlighted the advantages that MR offers as a ubiquitous tool for work. For instance, MR’s ability to personalize a user’s visual work environment has been shown to reduce distractions and stress while inducing a state of flow [79, 51, 87]. Schneider et al. [81] and Grubert et al. [37] examined the potential privacy benefits of the HMD form factor, which restricts the visibility of content to the wearer. Last but not least, MR relieves information displays of physical limitations [11, 37]. Using an HMD to present virtual content, including conventional windows, icons, and menus, MR provides more flexibility in where and how information is displayed [70]. As promoted by providers like Apple, Meta, and Sightful, it effectively provides a boundless workspace. This expanded screen real-estate may facilitate complex multi-window operations by serving as a type of external memory [72, 2, 8], enable more efficient multitasking [24], provide quicker peripheral access to information [58, 25, 38], and enhance immersion [27].
@@ 2.2 Window-based Mixed Reality
There has also been a persistent interest in using MR devices to enhance knowledge work by expanding traditional 2D displays and windows beyond the limitations of physical screens and into the third dimension [72]. Early work, such as Feiner et al. [32], explored various methods to register windows, including head-fixed and world-fixed approaches, as well as how to create connections between windows and physical objects. Raskar et al. [77] investigated the possibilities of expanding an office space with AR using projectors and cameras. Many recent commercial MR workspaces, such as the Varjo Workspace [89], Immersed [43], and vSpatial [90], now enable streaming of desktop elements into MR as virtual windows and support conventional input devices (i.e., mouse and keyboard) for interaction. Using conventional 2D displays and desktop input devices for MR can ease the transition to virtual displays for users coming from personal computers, while also providing access to familiar and highly refined interfaces and applications [72]. Therefore, in our study, we focus on emergent user behaviors when using an AR laptop that features a window-based immersive display along with keyboard and touchpad input.
Several prior studies have investigated the usage of window-based MR. For example, Ens et al. [28] classified window-based mixed reality workspaces in seven dimensions, such as whether the windows used an ego- or exocentric reference frame or were movable. Several prior studies [20, 70, 55] observed that users preferred to center main task content while using peripheral areas for supporting information. Su and Bailey [85] recommended against placing virtual contents behind the user. McGill et al. [62] reported on several common window configurations for general productivity tasks, including vertical, horizontal, two-plus-two, and following a “personal cockpit” [29]. As immersive technologies become more pervasively used, the context of use also becomes a critical consideration [36]. For example, the physical environment can hinder or enhance interactions [18]. Similarly, social acceptability has been shown to play an important role in the use of virtual monitors in airplanes [66] and other shared transit modalities [63]. While many usage patterns have been observed in shorter laboratory studies, our research investigates how these patterns may manifest in more realistic environments and over extended periods.
@@ 2.3 In-the-Wild and Longitudinal Studies of MR
There is a growing number of in-the-wild and extended-use studies of MR systems. In-the-wild studies of MR have demonstrated effectiveness across various application areas, including education [75], accessibility [80], military training [49], and on-the-go information access [15, 57]. Researchers have also investigated the long-term effects of MR usage on depth perception [48], social interactions [41, 6], and within industrial scenarios [35]. The most relevant to our work are previous studies on the use of MR for everyday productivity tasks.
In early investigations, many researchers engaged in informal self-experimentation. Starner [84] and Mann [14] independently designed and wore AR headsets almost daily for over a decade. Recent studies have continued this line of work with more participants, often with the aim of quantifying the effects. For example, Steinicke and Bruder [83] and Nordahl et al. [68] examined the impact of prolonged VR usage on key usability factors, including simulator sickness, over periods of 24 hours and 12 hours, respectively. Lu et al. [59] evaluated a glanceable AR prototype with three participants over three days. Guo et al. [39] and Shen et al. [82] report on a study in which 27 participants worked eight hours each in virtual and physical office environments, focusing on emotional and physiological needs, as well as visual and physical discomfort. Biener et al. [11, 9] compared VR work with a regular physical environment over five days with 16 participants. Pavanatto et al. [70] recently examined user and bystander experiences of Extended Reality (XR) usage in public, studying how users perform their own tasks in XR at various campus locations (e.g., library, dining). Complementing these empirical studies, Tran et al. [88] observed a keen interest in using MR for productivity through an ethnography of YouTube videos.
@@ 3.1 Apparatus
After pilot testing several devices and approaches, we chose to use the Sightful Spacetop EA (Early Access), shown in Figure 1. This standalone device consists of a pair of optical see-through display glasses (Xreal Light, $85$ grams, OLED display panels, $1920\times 1080$ pixels per eye, $72$ Hz refresh rate, $52^{\circ}$ diagonal field of view) for output attached to a standard keyboard and trackpad for input. It has a Qualcomm Snapdragon XR2 processor, a Kryo 585TM 8-core 64-bit CPU, an AdrenoTM 650 GPU, and 8GB of LPDDR5 RAM. The device runs a custom operating system, Spacetop OS, which enables users to flexibly open, close, resize, and move one or more window-based web applications (e.g., Google workspace application, WhatsApp, ChatGPT) on top of a virtual “spatial canvas” — a unified display space that is curved over a cylinder surrounding the user, similar to the concept described by Pavanatto et al. [72]. The spatial canvas is notably world-anchored [32] and can be zoomed in and out of, panned from side to side, height adjusted, tilted, and re-centered over the device keyboard. Participants could adjust the distance and size of windows and fonts to be readable comfortably. We qualitatively confirmed that at a 2-meter distance between the user and the virtual windows, a font size of 12 pt could be read comfortably, with visible degradation of legibility for fonts smaller than 8 pt. Legibility was sufficient for everyday productivity tasks, e.g., working in browser windows. The device does not have a front-facing camera, and does not have the capability to track the environment or other devices for spatial anchoring. The device’s technical specifications and supported interactions are described in detail in Supplementary Material. We did not make study-specific customizations to the device software for our deployment. We chose the device over video passthrough options like the Meta Quest 3 [64] or Apple Vision Pro [3] mainly because of its lighter and more ergonomic glasses form factor, which is essential for prolonged productive tasks [11].
@@ 3.2 Experimental Design and Procedure
Our particular approach was guided by several studies of methodological similarity (e.g., [23, 13, 44]). Diary studies are a common qualitative research method in HCI [22], widely used to understand behavior over longer periods, such as working within multi-device ecologies [44] or work-from-home conditions during the COVID-19 pandemic [13]. The study was approved by the Institutional Review Board (IRB) of the participating research institutions.
@@
We deployed the device to each participant for two weeks (14 days). During the usage period, participants were instructed to complete at least one 30-minute session on 10 of the 14 study days. Participants could complete as many additional sessions as they wished for extra compensation. This means they used the device during at least 10 different days.
Participants were encouraged to use the device during their daily activities and tasks, which they would typically complete regardless of their participation in the study, in the familiar locations they frequented as part of their routines. We deliberately chose not to set specific constraints on the task and environment, in order to gather more ecologically valid data on participants’ usage behaviors. We set individual sessions at a minimum of 30 minutes to ensure that participants had enough time to engage in meaningful tasks. We evaluated usage across 10 sessions to allow participants to acclimate to the device and understand how their usage changes with different tasks and extended exposure, without overburdening them.
@@
We conducted pilot tests of the survey with four knowledge workers, who were recruited through snowball sampling. 1 pilot participant completed 2 sessions (0.9 hours of total usage time), 2 completed six each (3.7 and 5.0 hours), and 1 completed 7 (4.9 hours). These pilot sessions helped us identify core questions, clarify wording, and ensure data quality. We revised the survey after each round of feedback. The results from these pilot sessions were not included in our analysis.
@@
Participants received $\mathdollar 150$ for completing ten usage sessions and the exit interview. They received $\mathdollar 15$ for each additional session they completed beyond the minimum requirements. Usage sessions took place from March 19, 2024, to August 29, 2024. We deployed between two and three devices simultaneously to speed up data collection.
@@ 3.3 Participants
We recruited 14 people (aged 22 to 56 years, $M$ = 30, $SD$ = 11; 9 male, 5 female) with normal or corrected vision to participate in the study. We recruited university students, researchers, and staff for our study due to their diverse tasks, extensive computer use in their work, and availability. Participants self-reported low to moderate experience with AR ($M$ = 2.6, $SD$ = 1.8) and VR ($M$ = 3.1, $SD$ = 1.7) on a scale from 1 (low) to 7 (high). Participants indicated that they regularly use 2 to 7 ($M$ = 4, $SD$ = 1) digital devices (i.e., mobile phones, laptops, tablets, and smartwatches). They reported spending approximately 5 to 14 hours each day ($M$ = 7, $SD$ = 2) on their primary device, which was either a laptop or desktop computer. All participants reported working remotely to some extent, typically at least once a week. They also reported having normal or corrected-to-normal vision.
@@ 3.4 Data Collection and Analysis
In addition, participants collectively submitted 140 photographs of their virtual workspaces and 109 photographs of their physical environments. All interviews with participants were automatically transcribed and then reviewed and corrected by the first author.
Figure 2: Overview of the themes that surfaced during our analysis. We report on how considerations, such as the participants’ tasks, influenced their workspace arrangement behaviors, such as the number of windows they opened. In addition, we highlight alternative workflows that involve the use of AR in tandem with physical displays and tasks. Finally, we summarize participants’ perceptions of the value and challenges of using AR.
@@
The final codebook comprised 171 individual codes (see Supplementary Material) that detailed the tasks participants performed using the AR laptop, the physical environments in which they worked, how they organized their virtual workspace for their tasks, and their perceptions regarding the value and limitations of the device. Using this shared codebook, two authors collaboratively conducted a final round of coding on the original data. They applied the codes from the codebook to each instance, labeling every survey entry and interview with zero or more codes, based on a unanimous decision. We do not report inter-rater reliability because following best practices from Hammer and Berland [40], coding for each instance was unanimously agreed upon by annotators. The same two authors also led the clustering process, organizing the codes into themes. These themes were refined iteratively with feedback from the remaining members of the research team, resulting in the findings presented below.